OpenStudy (anonymous):

differentiate the function ln(x(sqr root x^2-1))

OpenStudy (anonymous):

\[ \ln(x\sqrt{x^2-1}) = \ln(x) + \ln(\sqrt{x^2-1})=\ln(x) + \frac{1}{2}\ln(x^2-1) \]

OpenStudy (anonymous):

Then use chain rule.

OpenStudy (anonymous):

Or without chain rule: \[ \ln(x) + \frac{1}{2}\ln(x^2-1) = \ln(x)+\frac{1}{2}\ln[(x-1)(x+1)] = \ln(x)+\frac{1}{2}\ln(x-1)+\frac{1}{2}\ln(x+1) \]

OpenStudy (anonymous):

\[ \ln(x)+\frac{1}{2}\ln(x-1)+\frac{1}{2}\ln(x+1) \]

OpenStudy (anonymous):

thanx

hartnn (hartnn):

just a note, differentiating ln (x-1) is still a chain rule.

OpenStudy (anonymous):

thanx